OpenStudy (anonymous):

WILL MEDAL!! Order each group of quadratic functions from widest to narrowest graph. 16. y=3x^2, y=2x^2, y=4x^2 17. f(x)=5x^2, f(x)=-3x^2, f(x)=x^2 18. y=-1/2x^2, y=5x^2, y=-1/4x^2 19. f(x)=-2x^2, f(x)=-2/3x^2, f(x)=-4x^

OpenStudy (anonymous):

16. y=2x^2, y=3x^2, y=4x^2 17. y=-3x^2, y=x^2, y=5x^2 18. y=-1/4x^2, y=-1/2x^2, y=5x^2 19. y=-2/3x^2, y=-2x^2, y=-4x^2